two identical particles, each with a mass of 2.0 mg and a charge of 25 nC, are released simultaneously from rest when the two are 4.0 cm apart. what is the speed of either particle at the instant when the two are separated by 10 cm?
Answer #1

We use the conservation of energy:

E_o = E_f

U_o +K_o = U_f+K_f

\frac{kq^2}{d_1}+0 = \frac{kq^2}{d_2^2}+2\left ( \frac{1}{2}mv^2 \right )

then using given data:

\frac{9*10^9(25*10^{-9})^2}{0.04}+0 = \frac{9*10^9(25*10^{-9})^2}{0.1}+ 2\left ( \frac{1}{2}(2*10^{-6})v^2 \right )

solving for v

v = 6.50 m/s